Find p and q so that y = 2 x is a tangent, line at x = 3, to the curve
y = p x 2 + q x + 2

Grade:12

1 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear nitin,

  • The point of tangency is on the tangent line y = 2x, hence its y coordinate can also be calculated and its coordinates given by

    (3 , 2(3)) = (3 , 6)

  • The above point of tangency is also on the curve
    y = p x 2 + q x + 2, hence


    6 = p(3) 2 + q (3) + 2

  • Let us now find y '

    y ' = 2 p x + q

  • At x = 3 y ' is equal to the slope of the tangent line y = 2 x. Hence

    2 = 2 p (3) + q

  • We now solve the two linear equations

    9 p + 3 q = 4 and 6 p + q = 2

  • to find 9 p + 3 q = 4 and 6 p + q = 2

    p = 2 / 9 and q = 2 / 3

We are all IITians and here to help you in your IIT JEE preparation.

All the best.

 If you like this answer please approve it....

win exciting gifts by answering the questions on Discussion Forum

 

Sagar Singh

B.Tech IIT Delhi